A 25.0 mL sample of 0.150 M HIO3 is titrated with 0.240 M Sr(OH)2.

Write the net ionic equation for the above titration reaction.
1
Expert's answer
2020-05-24T02:35:53-0400

Solution:

Molecular Equation:

2HIO3(aq) + Sr(OH)2(aq) = Sr(IO3)2(aq) + 2H2O(l)

Full Ionic Equation:

2H+(aq) + 2IO3-(aq) + Sr2+(aq) + 2OH-(aq) = Sr2+(aq) + 2IO3-(aq) + 2H2O(l)

Net Ionic Equation:

2H+(aq) + 2OH-(aq) = 2H2O(l)

H+(aq) + OH-(aq) = H2O(l)


Answer:

Net Ionic Equation: H+(aq) + OH-(aq) = H2O(l)